The vectors field are specified as:


F = −12ax +5ay −10az

and G = 6ax +5ay −2az

Determine:

(a) The unit vector in the direction of −F + 3G

(b) The magnitude of 5ax + G − 2F

(c) |2F| |3G| (F+ G)

(d) The cross product between F and G


1
Expert's answer
2021-07-07T11:42:41-0400

Gives

"F=-12a_x+5a_y-10a_z"

"G=6a_x+5a_y-2a_z"

Part(a)

-F+3G=

"-(12a_x+5a_y-10a_z)+3(6a_x+5a_y-2a_z)"

"-F+3G=6a_x+0a_y+4a_z"


"P=-F+3G"

"\hat{P}=\frac{P}{|P|}=\frac{1}{\sqrt52}(6a_x+4a_z)"

Part (b)


"5a_x+G-2F=5a_x+6a_x+5a_y-2a_z-2(-12a_x+5a_y-10a_z)=35a_x-4a_y+18a_z"

"|5a_x+G-2F|=\sqrt{35^2+4^2+18^2}=\sqrt{1553}=39.40"

Part(c)

2F="-24a_x+10a_y-20a_z"

"3G=18a_x+15a_y-6a_z"

"F+G=-6a_x+10a_y-12a_z"


"|2F| |3G|(F+G)=\sqrt{1076}\times\sqrt{585}\times\sqrt{280}=13275.8"

Part(d)

"F\times G=\begin{bmatrix}
 a_x&a_y&a_z\\
 -12&5&-10\\
6&5&-2
\end{bmatrix}"


"F\times G=a_x(-10+50)-a_y(24+60)+a_z(-60-30)"

"F\times G=-40a_x-84a_y-90a_z"